Solve the following inequality: -5(3-x)less than or equal to -9x -8
Answer by checkley79(3341)

-5(3-x)<=-9x-8
-15+5X<=-9X-8
5X+9X<=-8+15
14X<=7
X<=7/14
X<=1/2 OR 0.5 ANS.
PROOF:
-5(3-.5)<=-9*.5-8
-5*2.5<=-4.5-8
-12.5<=-12.5